# An explicit lower bound for large gaps between some consecutive primes

### v3

@@ Abstract
Ford, Maynard and Tao [6] proved that there exists an effective abosolute constant $c_{LG}>0$ such that
@@
holds for any sufficiently large $X$. The main purpose of this paper is to determine the constant $c_{LG}$ above. We see that $c_{LG}$ is determined by several factors related to analytic number theory, for example, the ratio of integrals of functions in the multidimensional sieve of Maynard [14], the distribution of primes in arithmetic progressions to large moduli, and the coefficient of upper bound sieve of Selberg. We prove that the above inequality is valid at least for $c_{LG}\approx 2.0\times 10^{-17}$.
@@ 1 Introduction
and in particular, put $G(X):=G_{1}(X)$, which denotes the largest prime gap below $X$. The prime number theorem yields the number of primes below $x$ is asymptotically $x/\log x$, so the average gap of primes below $x$ is $(1+o(1))\log x$. Hence we get the trivial lower bound
@@
Cramér constructed a probabilistic model of the set of prime numbers and conjectured that $G(X)\sim(\log X)^{2}$. (See also refinements of Cramér’s model by Granville [8], Firoozbakht [20] and Wolf [23].) In 1931, Westzynthius [22] made the first quantitative improvement on the trivial bound and proved
@@
where $\log_{n}$ denotes the $n$ times composition of the logarithmic function, i.e., $\log_{n}X:=(\underbrace{\log\circ\cdots\circ\log}_{n})X$. In particular, it follows that $G(X)/\log X\to\infty$ as $X\to\infty$, so $G(X)$ can be arbitrarily large compared with the average gap. Erdős [2] improved on Westzynthius’ result and obtained
@@
holds with $c=\frac{1}{3}$. Several mathematicians improved on the value of the coefficient $c$ above (see Schönhage [21] with $c=\frac{1}{2}e^{\gamma}$, Rankin [19] with $c=e^{\gamma}$, Maier and Pomerance [13] with $c=1.31256e^{\gamma}$, and Pintz [16] with $c=2e^{\gamma}$). Erdős conjectured that (1.1) holds with arbitrarily large $c$. This problem had been unsolved for a long time, but a team of Ford, Green, Konyagin, Tao [5] and Maynard [14] independently solved in August of 2014. Several months later, these five mathematicians [4] worked together and proved
$G(X)\gg\frac{\log X\log_{2}X\log_{4}X}{\log_{3}X},$ (1.2)
which was the first quantitative improvement of Rankin’s bound (1.1) in almost 80 years.
@@
as an application of his famous Maier matrix method and Pintz’s ideas in [16], where $(X_{n})$ is some monotonically increasing sequence such that $X_{n}\to\infty$ as $n\to\infty$. Pintz [17] improved on Maier’s result and proved
@@
An issue of Maier’s argument is that one has to avoid Siegel zeros, and this is the reason why the results of Maier and Pintz above are restricted to a special sequence $(X_{n})$, rather than all sufficiently large $X$.
@@
holds as $X\to\infty$, and the implied constant above is absolute and effective. In other words, for any sufficiently large $X\geq 3$, there exists an absolute computable constant $c_{LG}$ such that
@@
holds. In addition to the technique to handle the possibility of the existence of exceptional zeros of $L$-functions, their result relies on hypergraph covering theorem, the construction of sieve weight, and very clever and highly technical probabilistic arguments in [4].
@@
Of course, this value is not the best that can be obtained with current techniques and could be improved to some extent with some effort. In any case, however, as long as we rely on current methods, the coefficient is likely to have to be fairly small.
@@ 2 Notation
Let $10<\theta<$ 1 be a parameter of Hypothesis 1 of [14]. Note that in our situation, unconditionally one can take
@@ 4 Possible values of $C_{UB}$ and $D_{UB}$ in Assumption UB
provided that $\log x=o(\log Z)$. Since $Z\geq P^{D_{UB}}=e^{(1+o(1))D_{UB}x}$, this condition is satisfies for any $D_{UB}>0$. Consequently, the Assumption UB is valid for
@@ 5 Possible values of $C_{PAP}$ and $D_{PAP}$ in Assumption PAP
holds uniformly for $x/Q\leq h\leq x$, $\exp(x^{\frac{1}{2}}\log x)\leq Q\leq x^{b}$. We need to make clear the value of $a$. Due to the above consideration, the $L$-functions associated to primitive characters modulo at most $T$ does not have zero in
@@
By Theorem 6 of [7], for some constant $c_{ZD}>0$,
@@
holds uniformly for $0\leq\alpha\leq 1$. Here, $N(\alpha,T,\chi)$ denotes the number of zeros of $L(s,\chi)$ in the domain $\alpha\leq\Re(s)\leq 1$, $|\Im(s)|\leq T$. With this $c_{ZD}$, Theorem 7 of [7] is valid for $T^{c_{ZD}}\leq x^{\frac{1}{2}}$, $T=Q^{5}$. Therefore, for $\exp(\log^{\frac{1}{2}}x)\leq Q\leq x^{\frac{1}{10c_{ZD}}}$, (5.2) holds for $\frac{x}{Q}\leq h\leq x$. Therefore, the proof og Lemma 2 of Maier [12] is valid for $x\geq Q^{10c_{ZD}}$.
In [12], the contribution of the terms with $\chi=\chi_{0}$ to $\psi(x;q,a)$ is $(1+o(1))x/\varphi(q)$, whereas that of the terms with $\chi\neq\chi_{0}$ is $O_{\leq}(xe^{-aD}/\varphi(q))$. Therefore, we have
@@
Let $x$ be a sufficiently large real number, $B_{0}\in\mathbb{N}$ and put
@@
(Proof of Proposition 6.2) Proposition 6.2 is obtained by applying the sieve of Maynard (Theorem 6 of [6]. For the proof, see Section 6 of [4]). In particular, the constant $c_{0}$ in (6.1) of [6] can be chosen as $c_{0}=\frac{1}{5}$, following the argument of Maynard [14]. The $C$ in Theorem 5 of [6] is computed in the following way. Put

### v4

@@ Abstract
Ford, Maynard and Tao [6] proved that there exists an effective absolute constant $c_{LG}>0$ such that
@@
holds for any sufficiently large $X$. The main purpose of this paper is to clarify the numerical value of the constant $c_{LG}$ such that the above inequality holds. We see that $c_{LG}$ is determined by several factors related to analytic number theory, for example, the ratio of integrals of functions in the multidimensional sieve of Maynard [14], the distribution of primes in arithmetic progressions to large moduli, and the coefficient of upper bound sieve of Selberg. We prove that the above inequality is valid at least for $c_{LG}\approx 2.0\times 10^{-17}$.
@@ 1 Introduction
and in particular, put $G(X):=G_{1}(X)$, which denotes the largest prime gap below $X$. The prime number theorem yields the number of primes below $x$ is asymptotically $x/\log x$, so the average gap of consecutive primes below $x$ is $(1+o(1))\log x$. Hence we get the trivial lower bound
@@
Harald Cramér [1] constructed a simple probabilistic model of the set of prime numbers, and following his model, he conjectured that $G(X)\sim(\log X)^{2}$. (See also refinements of Cramér’s model by Granville [8], Firoozbakht [20] and Wolf [24].) In 1931, Westzynthius [23] made the first quantitative improvement on the trivial bound and proved
@@
where $\log_{n}$ denotes the $n$ times composition of the logarithmic function, i.e., $\log_{n}X:=(\underbrace{\log\circ\cdots\circ\log}_{n})X$. In particular, it follows that $G(X)/\log X\to\infty$ as $X\to\infty$, so $G(X)$ can be arbitrarily large compared with the average gap. In 1934, Ricci [21] slightly improved this and showed $G(X)\gg\log X\log_{3}X$. Erdős [2] improved on Westzynthius’ result and obtained
@@
holds with $c=\frac{1}{3}$. Several mathematicians improved on the value of the coefficient $c$ above (see Schönhage [22] with $c=\frac{1}{2}e^{\gamma}$, Rankin [19] with $c=e^{\gamma}$, Maier and Pomerance [13] with $c=1.31256e^{\gamma}$, and Pintz [16] with $c=2e^{\gamma}$). Erdős conjectured that (1.1) holds with arbitrarily large $c$. This problem had been unsolved for a long time, but Maynard [14] and a team of Ford, Green, Konyagin, Tao [5] independently solved at almost the same time in August 2014. Several months later, in a joint work [4], they obtained
$G(X)\gg\frac{\log X\log_{2}X\log_{4}X}{\log_{3}X}.$ (1.2)
This was the first quantitative improvement of Rankin’s bound (1.1) in almost 80 years.
@@
by combining his famous Maier matrix method and Pintz’s ideas in [16], where $(X_{n})$ is some monotonically increasing sequence such that $X_{n}\to\infty$ as $n\to\infty$. Pintz [17] improved on Maier’s result and proved
@@
An issue of Maier’s argument is that one has to avoid (possible) Siegel zeros, and this is the reason why the results of Maier and Pintz above are restricted to a special sequence $(X_{n})$, rather than all sufficiently large $X$.
@@
holds as $X\to\infty$, and the implied constant above is absolute and effective. In other words, for any sufficiently large $X\geq 3$, there exists an absolute computable constant $c_{LG}>0$ such that
@@
holds. In addition to the technique to handle the possibility of the existence of exceptional zeros of $L$-functions, their result relies on hypergraph covering theorem, the construction of multidimensional sieve weight, and very clever and highly technical probabilistic arguments in [4].
@@
Obviously, this value is not the best that can be obtained with current techniques and could be improved to some extent with some effort, for example, by applying recent results on zero density estimates and explicit zero free regions of Dirichlet $L$-functions. In any case, however, as long as we rely on current methods, the coefficient is likely to have to be fairly small.
@@ 2 Notation
Let $10<\theta<$ 1 be a parameter of Hypothesis 1 of [14]. In our situation, unconditionally one can take
@@ 4 Possible values of $C_{UB}$ and $D_{UB}$ in Assumption UB
provided that $\log x=o(\log Z)$. Since $Z\geq P^{D_{UB}}=e^{(1+o(1))D_{UB}x}$, this condition is valid for any $D_{UB}>0$. Consequently, the Assumption UB is valid for
@@ 5 Possible values of $C_{PAP}$ and $D_{PAP}$ in Assumption PAP
holds uniformly for $x/Q\leq h\leq x$, $\exp(x^{\frac{1}{2}}\log x)\leq Q\leq x^{b}$. We need to make clear the value of $a$. Due to the above consideration, the $L$-functions associated to primitive characters modulo at most $T$ do not have zero in
@@
Next, by Theorem 6 of [7], for some constant $c_{ZD}>0$,
@@
holds uniformly for $0\leq\alpha\leq 1$. Here, $N(\alpha,T,\chi)$ denotes the number of zeros of $L(s,\chi)$ in the domain $\alpha\leq\Re(s)\leq 1$, $|\Im(s)|\leq T$. With this $c_{ZD}$, Theorem 7 of [7] is valid for $T^{c_{ZD}}\leq x^{\frac{1}{2}}$, $T=Q^{5}$. Therefore, for $\exp(\log^{\frac{1}{2}}x)\leq Q\leq x^{\frac{1}{10c_{ZD}}}$, (5.2) holds for $x/Q\leq h\leq x$. Therefore, the proof og Lemma 2 of Maier [12] is valid for $x\geq Q^{10c_{ZD}}$.
In [12], the contribution of the terms with $\chi=\chi_{0}$ to $\psi(x;q,a)$ is $(1+o(1))x/\varphi(q)$, whereas that of the terms with $\chi\neq\chi_{0}$ is $O_{\leq}(xe^{-aD}/\varphi(q))$. Here, $f(x)=O_{\leq}(g(x))$ means that $|f(x)|\leq g(x)$ holds for any sufficiently large $x$. Therefore, we have
@@
Let $x$ be a sufficiently large real number and put
@@
(Proof of Proposition 6.2) Proposition 6.2 can be obtained by applying the sieve of Maynard (Theorem 6 of [6]. For the proof, see Section 6 of [4]). In particular, the constant $c_{0}$ in (6.1) of [6] can be chosen as $c_{0}=\frac{1}{5}$, following the argument of Maynard [14]. The quantity $C$ in Theorem 5 of [6] is computed in the following way. Put